OpenStudy (anonymous):

y^4/5=34 Help please!

OpenStudy (anonymous):

Use SADMEP. So divide first by 5 on both sides.

OpenStudy (anonymous):

i mean multiply

OpenStudy (anonymous):

So would it be (5)y^4/5(5)?

OpenStudy (anonymous):

No, you multiply on each side. On the left side it would cancel out the /5 and on the right side you would have 34*5

OpenStudy (anonymous):

Got cha

OpenStudy (anonymous):

good

OpenStudy (anonymous):

Do you know what to do next?

OpenStudy (anonymous):

add?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

then what do I do next??

OpenStudy (anonymous):

first of all what is 34*5?

OpenStudy (anonymous):

170

OpenStudy (anonymous):

so now the equation reads y^4=170

OpenStudy (anonymous):

Alright then from here what's next?

OpenStudy (anonymous):

you have to put both sides to the fourth root. If you have a calculator it be good to use it...

OpenStudy (anonymous):

Okay let me calculate it and then tell me if I'm right.

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

170 rooted is 13.03?

OpenStudy (anonymous):

Wait is the question (y^4)/5=34 or (y^(4/5)=54?

OpenStudy (anonymous):

its y with 4/5 as the exponent equals 34?

OpenStudy (anonymous):

so I'm solving for y.

OpenStudy (anonymous):

Ohhhh okay, one second then...

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

the closest im getting is 82^(4/5)=33.967 Do you need more specific?

OpenStudy (anonymous):

Nope that will work thank u so much:)

OpenStudy (anonymous):

Most exact is 82.10103769

OpenStudy (anonymous):

Can I get a Medal?

OpenStudy (anonymous):

Thanks && Of course :)

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

No problem